Have I correctly solved the 4 problems below using the systems of simultaneous equations, using either the substitution method or the addition method?
Problem 1
x+2y=15
x-2y=-9
2x =6
x=3
3+2y=15
2y=15-3
2y=12
y=6
Problem 2:
x-y=0
7x-3y=24
-3(x-y)=0
-3x+3y=0
7x-3y=24
4x=24
x=6
6-y=0
-y=-6
y=6
Problem 3:
3x-8y=10
x-4y=3
-2(x-4y)=-2(3)
-2x+8y=-6
3x-8y=10
x=4
3(4)-8y=10
12-8y=10
-8y=10-12
-8y=-2
y=1/4
Problem 4:
2x+3y=-5
3x-y=20
3(3x-y)=3(20)
9y-3y=60
9y-3y=60
2x+3y=-5
11x=55
x=5
2(5)+3y=-5
10+3y=-5
3y=-5-10
3y=-15
y=-5
Thank you

Answer by ikleyn(52788)

.

Have I correctly solved the 4 problems below using the systems of simultaneous equations, using either the substitution method or the addition method?

Problem 1

x + 2y = 15
x - 2y = -9

2x = 6

x = 3

3 + 2y = 15
2y = 15-3
2y = 12

y=6



Problem 2:

x - y = 0
7x - 3y = 24

-3(x - y) = 0

-3x + 3y = 0
7x - 3y = 24

4x = 24

x = 6

6 - y = 0
-y = -6

y = 6



Problem 3:

3x - 8y = 10
x - 4y = 3

-2(x - 4y) = -2(3)

-2x + 8y = -6
3x - 8y = 10

x = 4

3(4) - 8y = 10
12 - 8y = 10
-8y = 10-12
-8y = -2

y = 1/4



Problem 4:

2x + 3y = -5
3x - y = 20

3(3x - y) = 3(20)
9y - 3y = 60

9y - 3y = 60
2x + 3y = -5

11x = 55

x = 5

2(5) + 3y = -5
10 + 3y = -5
3y = -5-10
3y = -15

y = -5

Thank you

~~~~~~~~~~~~~~~~~


Hello,  I checked every line in your post.

They  ALL  are correct.

Very good job.  My congratulations  (!)


Now,  there is another method to check,  instead of checking every line.

After completing the solution,  you can / (you should) / (you MUST)  substitute the obtained numbers  (the answer)
into the original equations  TO  ASSURE  that you get valid equalities.

Such checking is  INDISPENSABLE  and  NECESSARY  part of the solution.
It is assumed that every student  KNOWS  about it an  MAKES  IT  after solving each and every such problem.


Well done (!)